OpenStudy (anonymous):

What is the vertex form of the equation y=-x^2+12x-4 I am looking for help. I need to know step by step how to do it.

OpenStudy (anonymous):

@bibby can u help

OpenStudy (bibby):

can you complete the square

OpenStudy (bibby):

@Crazymoron

OpenStudy (anonymous):

are you in connections academy?

OpenStudy (anonymous):

im stuck on this problem to so i can not help sorry

OpenStudy (bibby):

I figured I'd try helping the both of you in one problem

OpenStudy (anonymous):

y = x^2 + 12x - 4 y = x^2 + 12x + (36 - 36) - 4 y = (x^2 + 12x + 36) - 36 - 4 y = (x + 6)^2 - 40 @Crazymoron

OpenStudy (bibby):

\(y=-x^2-12x+4\\y=-1(x^2+12x-4)\\y+36=-1(x^2+12x+36-4)\\y+36=-1((x+6)^2-4)\\y+36=-(x+6)^2+4\\y=-(x+6)^2-32\)

OpenStudy (bibby):

is this wrong?

OpenStudy (anonymous):

yeah the 4 is negative in the equation

OpenStudy (anonymous):

i did it totally different. i used method 2 on this website http://mathbitsnotebook.com/Algebra1/Quadratics/QDVertexForm.html so i ended up geting Y=-x(x-6)^2-40

OpenStudy (bibby):

the x^2 is negative

OpenStudy (anonymous):

i know. i did everything the same

OpenStudy (anonymous):

as it was on the website

OpenStudy (anonymous):

@kimbllse777

OpenStudy (anonymous):

this is how I solved it.... Hope it is right...